geerky42 (geerky42):

In an amusement-park ride, riders stand with their backs against the wall of a spinning vertical cylinder. The floor falls away and the riders are held up by friction. The acceleration of gravity is 9.81 m/s². If the radius of the cylinder is 10 m, find the minimum number of revolutions per minute necessary when the coefficient of static friction between a rider and the wall is 0.5. Answer in units of rev/min.

geerky42 (geerky42):

@lachlanrob @Algebraic! @NoelGreco

OpenStudy (anonymous):

I think the question is asking you to calculate the minimum number of rpm to keep the riders from falling. I don't know the equation for this though sorry.

OpenStudy (anonymous):

the normal force between the wall and the rider provides the centripetal force. u*N must be => mg N=mg/u mg/u =mV^2/r sqrt(rg/u) = V you can find RPMs easily now?

geerky42 (geerky42):

could you clarify the question by draw it out? i know what question is asking for but i couldn't understand exactly what is going on... i have bad reading comprehension...

OpenStudy (anonymous):